OpenStudy (anonymous):

solve the quation: 7/x+1=5/x-3

OpenStudy (anonymous):

on solving 2/x=-4 -1/2

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

x=-1/2

OpenStudy (anonymous):

i dont have that as a choice

OpenStudy (anonymous):

-8,8,11,13

OpenStudy (anonymous):

but it is the answer if u still have doubt put x=-1/2 & the equation will hold

OpenStudy (anonymous):

okay thank you. i dont have a doubt but all this stuff is on the computer

OpenStudy (anonymous):

you better write the question with brackets may be that's the problem

OpenStudy (anonymous):

I dont know. it just gave it to me like i wrote it out

OpenStudy (anonymous):

I got it answer is 13

OpenStudy (anonymous):

thank you so much.